OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form.

OpenStudy (anonymous):

@Aditya790 Can You Help Me Please?

OpenStudy (anonymous):

The slope of required line = 1, which is parallel to line y = -x + 6 As the lone passes through (-4,2) It can be written as (y-2)= 1.(x+4) The required line is y = x + 6

OpenStudy (anonymous):

Is This Correct?

OpenStudy (anonymous):

parallel lines have the same slopes y = -x + 6 has a slope of -1 so the equation of the line you are looking for has to have the slope of -1 and pass through -4,2 y = -1x + b Substitute x = -4, y = 2 to find b, and there you got the equation for the line.